weather balloon

A weather balloon contains 233 L ofhelium at 22 C and 760. mm Hg. What is the volume of the balloon when it ascends to an altitude where the temperature is-54 C and 511 mm Hg?
Answer #1

considering PV/T = constant

We get , P1V1/T1 = P2V2 /T2 , V1 = 233 L , T1 =273 +22 = 295 K , T2 = 273 -54 = 219 K , P1 = 760 mm ,P2 = 511 mm

So, V2 = P1V1T2/(T1P2) = 760 x 233 x 219 / (295 x 511) = 257.3 L